# CATNet: A geometric deep learning approach for CAT bond spread prediction in the primary market

### v1

@@ 5.3.1 Core risk metrics
5.3.2 Contractual and market-based beatures
@@ 5.3.3 Network topology features
A novel contribution of this work is the R-GCN’s ability to exploit the graph structure of the CAT bond market, learning from the web of relationships connecting issuers, underwriters, and perils. The high predictive power of topological features confirms that an entity’s position and influence within this network carry significant pricing information. This finding provides a quantitative basis for the well-documented “issuer effects” and other dynamics in the CAT bond market (Chatoro et al.,, 2023).
@@ 6 Concluding remarks
Our results revealed a significant advantage of this approach. We found that the CAT bond market exhibits the characteristics of a scale-free network, a structure dominated by a few highly connected “hubs” and many sparsely connected entities. This architecture, while efficient, introduces systemic vulnerabilities, as shocks to these central nodes can propagate widely. CATNet’s success stems directly from its ability to navigate this structure. Its graph representation alone significantly outperformed a strong Random Forest benchmark, and the inclusion of topological features provided a further, substantial performance boost. More importantly, interpretability analysis showed that these centrality measures are not statistical artifacts; they are quantitative proxies for the influence of these network hubs, effectively capturing long-held industry intuition about issuer reputation, underwriter influence, and peril concentration.
@@ C.1 Network topology
Definition (Adjacency matrix). An adjacency matrix corresponding to a graph of size $N$ is a square matrix of size $(N\times N)$ whose $(i,j)$-th element, denoted as $A_{ij}$ [^12] For an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|V(G)|\times|V(G)|}$, we have that $\mathbf{A}[i,j]=A_{ij}$, is assigned a value of 1 if there exists a connection between node $i$ and $j$. Conversely, if there is no connection, the value is set to 0.
Definition (Undirected graph). A finite graph $G$ is called an undirected graph if all of its edges are undirected (i.e., lack a specific direction), for which the edge set is defined as $\mathcal{E}(G)=\{\{v_{i},v_{j}\}:v_{i},v_{j}\in\mathcal{V}(G)\}$ (for simplicity in notation, one can say $ij\in\mathcal{E}(G)$ in place of $\{v_{i},v_{j}\}\in\mathcal{E}(G)$). For the directed graph (also called digraph) the edge set is defined as $\mathcal{E}(G)=\{(v_{k},v_{u}),\dots,(v_{i},v_{j})\}$. Let $|\mathcal{E}(G)|$ and $|\mathcal{V}(G)|$ specify the sizes of the edge set and node set respectively, hence $|\mathcal{E}(G)|=L$ denotes the number of links and $|\mathcal{V}(G)|=N$ is the number of nodes in a graph. We add that to fit our purpose in this paper, we stick to a finite and undirected graph that is simple [^13] By a simple graph we mean that the graph does not have any multiple or self-loop links..
Definition (Complete graph). A finite graph $G$ of size $N$ is a complete graph (also called clique) if the edge set includes every possible pair of vertices, i.e., $\mathcal{E}(G)=\{\{v_{i},v_{j}\}:v_{i},v_{j}\in\mathcal{V}(G)\;\text{for}\;1\leq i\leq j\leq N\}$, that is, each node is connected to every other node.
Definition (Connected graph). A finite graph $G$ of size $N$ is a connected graph if and only if for every pair of distinct nodes $v_{i},v_{j}\in\mathcal{V}(G)$ there exists a path through which the two nodes are connected together. Hence, a complete graph is a fully connected graph.
Definition (Degree, degree distribution, and average degree). The degree of node $v_{i}\in\mathcal{V}(G)$ in the graph $G$ is defined as the number of links that node $i$ has with other nodes or equivalently the number of edges containing node $i$, that is, $d_{i}^{(G)}=\#\{v_{j}\in\mathcal{V}(G):\{v_{i},v_{j}\}\in\mathcal{E}(G)\}$. The collection of degrees of all vertices forms a sequence of degrees denoted by $\boldsymbol{d}=(d_{i}^{(G)})_{v_{i}\in\mathcal{V}(G)}$. The probability that a randomly selected node $o\in\mathcal{V}(E)$ has a degree equal to $k$ (shown by $D=d_{o}^{(G)}=k$, where $k=0,1,\dots$) is given by the following relation:
$p_{k}^{(G)}=\mathbb{P}(D=k)=\frac{1}{|\mathcal{V}(G)|}\sum_{v_{i}\in\mathcal{V}(G)}\mathbb{I}_{\{d_{i}^{(G)}=k\}}=\frac{N_{k}}{N}$ (C.1)
where $N_{k}$ is defined to be the number of nodes with degree $k$. The definition of degree for a node can be extended to the graph level, for which the average degree is computed as below:
$\langle G\rangle=\frac{1}{|\mathcal{V}(G)|}\sum_{v_{i}\in\mathcal{V}(G)}d_{i}^{(G)}=\frac{1}{|\mathcal{V}(G)|}\sum_{i=1}^{N}d_{i}^{(G)}=\frac{2L}{N}$ (C.2)
where $L=|\mathcal{E}(G)|$ is the total number of links in the graph. The range within which the average degree changes is between 0 (for an empty graph) and $|\mathcal{V}(G)|-1$ (for a complete graph). The other way to express relation (2) is:
$\mathbb{E}(D)=\sum_{k=0}^{+\infty}kp_{k}^{(G)}$ (C.3)
which shows the fact that the average degree of a graph can be governed by the degree distribution. Given that $p_{k}^{(G)}$ is the empirical probability mass function corresponding to the true underlying degree distribution, one can understand $\langle G\rangle$ as a realization of the true mean $\mathbb{E}(D)$.
Definition (Isolated nodes and hubs). Nodes with a degree of zero, denoted as $k=0$, are referred to as isolated nodes, while nodes with the highest number of connections in the graph are termed hubs. A takeaway point is that in real-world networks, there is often a significant disparity between the minimum and maximum degrees, resulting in a wide variation in node degrees.
Definition (Node’s Neighbors): The neighborhood of a given node in the graph is defined as all nodes directly connected to that specific node. Formally, let $G=(\mathcal{V}(G),\mathcal{E}(G))$ be a graph with $v_{i}\in\mathcal{V}(G)$, then the set of neighbors of node $v_{i}$ is defined to be $\mathcal{N}(v_{i})=\{v_{j}\in\mathcal{V}(E):\{v_{j},v_{i}\}\in\mathcal{E}(G)\}$.
Notably, for an undirected graph of size $N$, the adjacency matrix assumes symmetry, wherein $A_{ij}=A_{ji}$, and that $A_{ii}=0$ due to the lack of existing self-loops [^14] Note that in a weighted graph, $A_{ij}=w_{ij}$. The adjacency matrix corresponding to real-world networks typically exhibits sparsity, indicating that only a small portion of the matrix contains nonzero values. Essentially, the number of edges in a real network, $L=\frac{1}{2}\sum_{i,j}^{N}A_{ij}$, is considerably less than what would be found in a complete network of the same size, expressed as $L_{max}=\frac{N(N-1)}{2}$.
Definition (Path and Path’s length). A simple path is a route that starts at one node and ends at another, visiting each intermediate node no more than once. If the starting and ending nodes are the same, the path is referred to as a cycle. The length of the path is determined by the number of links it contains.
Definition (Graph distance). The graph distance between nodes $u$ and $v$, denoted by $dist_{G}(u,v)$, is defined to be the minimal number of edges in a path linking $u$ and $v$.
For instance, when $dist_{G}(u,v)=2$, it indicates that the shortest path between nodes $u$ and $v$ is of length 2, implying the existence of an intermediate node $l$ such that $A_{ul}A_{lv}=1$. Multiple two-step routes however may exist from node $u$ to node $v$. This insight highlights that the shortest path between nodes is not necessarily unique. Thus, it is possible to determine the number of shortest paths with a length of $2$ between nodes $u$ and $v$ through $N_{uv}^{(2)}=\sum_{l=1}^{N}A_{ul}A_{lv}$. It’s important to note that if nodes $u$ and $v$ are directly connected, then $dist_{G}(u,v)=1$.
Definition (Graph diameter). The graph diameter is defined to be the largest distance between any pair of nodes, denoted by $diam(G)=\max\limits_{u,v\in\mathcal{V}(G)}dist_{G}(u,v)$.
Definition (Average path). The average path which defines the average distance of all pairs in the graph is given by:
$\langle dist_{G}\rangle=\frac{1}{N(N-1)}\sum_{i,j=1,i\not=j}^{N}dist_{G}(i,j)$ (C.4)
Similar to what we discussed in the degree distribution, if $o_{1}$ and $o_{2}$ are two randomly drawn nodes from the node set $\mathcal{V}(G)$, where $G$ is assumed to be a connected graph (i.e., for any $u,v\in\mathcal{V}(G)$: $dist_{G}(u,v)<\infty$) and define $H=dis_{G}(o_{1},o_{2})$, then $H$ is a random variable which models all distances in the graph through the distance distribution below:
$\mathbb{P}(H=d)=\frac{\sum_{i,j=1,i\not=j}^{N}\mathbb{I}_{\{dist_{G}(i,j)=d\}}}{N(N-1)}$ (C.5)
Therefore, the expectation of $H$, i.e., $\mathbb{E}(H)$ can be understood as the average path length of the graph ($\langle dist_{G}\rangle$). Moreover, intuitively, the average path of a graph provides insight into the speed at which the information flows across the network, reflecting the graph’s efficiency. In practice, algorithms such as Breadth-First Search (BFS) and Depth-first Search (DFS) can be used to facilitate the task of finding the shortest path or detecting cycles within large graphs (see e.g., Labonne, (2023) and Stamile et al., (2021)).

### v2

@@ 5.3.1 Core risk metrics
5.3.2 Contractual and market-based features
@@ 5.3.3 Network topology features
As already discussed in Section 4.2.2, a novel contribution of this work is the R-GCN’s ability to exploit the graph structure of the CAT bond market, learning from the web of relationships connecting issuers, underwriters, and perils. The high predictive power of topological properties confirms that an entity’s position and influence within this network carry significant pricing information. This finding provides a quantitative basis for the well-documented “issuer effects” and other dynamics in the CAT bond market (Chatoro et al.,, 2023).
@@ 6 Concluding remarks
Our results revealed a significant advantage of this approach. We found that the CAT bond market exhibits the characteristics of a scale-free network, a structure dominated by a few highly connected “hubs” and many sparsely connected entities. This architecture, while efficient, introduces systemic vulnerabilities, as shocks to these central nodes can propagate widely. CATNet’s success stems directly from its ability to navigate this structure. CATNet significantly outperformed a strong Random Forest and XGBoost benchmarks. More importantly, interpretability analysis showed that these centrality measures are not statistical artifacts; they are quantitative proxies for the influence of these network hubs, effectively capturing long-held industry intuition about issuer reputation, underwriter influence, and peril concentration.
@@ B.1 Hyperparameters
B.2 Ablation study with macroeconomic variables
Table 11: OOS model performance across folds (with macro variables)
Fold Random Forest XGBoost R-GCN
MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%)
1 0.4021 64.03 0.3377 69.79 0.2813 74.83
2 0.1685 79.68 0.1983 76.08 0.1423 82.84
3 0.2879 75.33 0.2423 79.24 0.3394 70.92
4 0.2040 75.75 0.2302 72.63 0.2667 68.30
5 0.1260 79.75 0.1219 80.41 0.1478 76.25
6 0.4114 70.41 0.2479 82.18 0.2546 81.69
7 0.2039 82.59 0.2085 82.19 0.1911 83.68
8 0.3037 72.04 0.2939 72.94 0.3030 72.11
9 0.1672 77.98 0.1377 81.86 0.1638 78.43
10 0.2324 79.05 0.1808 83.70 0.2065 81.38
Average 0.2507 75.66 0.2199 78.10 0.2296 77.04
Table 12: OOT model performance across walk-forward splits (with macro variables)
Train / Val Split Test Year Random Forest XGBoost R-GCN
MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%)
1999–2014 — Val 2015 2016 0.2318 66.23 0.1994 70.95 0.2341 65.90
1999–2015 — Val 2016 2017 0.1299 78.09 0.1027 82.68 0.0526 91.12
1999–2016 — Val 2017 2018 0.3056 51.61 0.2662 57.86 0.1505 76.17
1999–2017 — Val 2018 2019 0.2411 67.57 0.2296 69.12 0.2869 61.41
1999–2018 — Val 2019 2020 0.3212 45.61 0.2905 50.81 0.2498 57.70
1999–2019 — Val 2020 2021 0.1509 78.14 0.1338 80.62 0.0458 93.36
Average 0.2301 64.54 0.2037 68.67 0.1700 74.28